OpenStudy (anonymous):

please help! What is the slope of a line that passes through the point (-2, 3) and is parallel to a line that passes through (3, 7) and (-2, -8)?

OpenStudy (anonymous):

=-8-7/-2-3 =-15/-3 =5

OpenStudy (anonymous):

\[{ (-8-7) \over (-2-3) }= 5\] This will be slope of the every parallel line, that is parallel to this line.

OpenStudy (anonymous):

the equation of the line form is: y-3=5(x+2) y-3=5x+10 5x-y+13=0

OpenStudy (anonymous):

I like you name haha_proof :D

OpenStudy (anonymous):

thanks..im not really good in math as you think..

OpenStudy (anonymous):

Thanks for all of your help :) and thank you :D

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

can you please help with one more im doing this exam and im so lost on these type of problems and i have to get a good grade.?

OpenStudy (anonymous):

you're welcome ...its our pleasure to solve math problems,hehe

OpenStudy (anonymous):

What is the slope of the line passing through the point (3, -4) and perpendicular to the line y = 3x + 5? Write your answer as a fraction.

OpenStudy (anonymous):

what you are going to do is that familiarize the different form of the line equation.

OpenStudy (anonymous):

y=3x+5 is in the form of slope intercept form where y=mx +b. so the slope is m=3. to get perpendicular slope of the given line you need to get the multiplicative inverse that will result to -1 so the slope is m=-1/3

OpenStudy (anonymous):

Thank you for all of your help :) im still a little confused but ill try my next problem on own!

OpenStudy (anonymous):

*my own

OpenStudy (anonymous):

ok,just remember that parallel lines do have the same slope and perpendicular lines get -1 if slopes are to be multiplied.

OpenStudy (anonymous):

ok thanks again!